OpenStudy (anonymous):

A hallway has an area of 63 square feet. The length of the hallway is 7 times longer than the width. What is the length of the hallway? @johnweldon1993

OpenStudy (anonymous):

Use substitution Lxw = area L = 7w 7w * w = 63 7w^2 = 63 w^2 = 9 w = 3 So the width is 3. Length is 7(3) = 21 Hope that helps.

OpenStudy (anonymous):

they want feet